### v1

# ADAPTIVE-$\lambda$ SISS in MACHINE UNLEARNING for DDPM & VAE

Abstract
Machine Unlearning (MU) is critical for large generative models, such as VAEs[1] and DDPMs[4], to enforce the right to be forgotten and ensure safe content generation without expensive retraining. Existing methods, like the Static-$\lambda$ SISS framework[6] for Diffusion Models, rely on a fixed data mixing parameter $\lambda$, which is suboptimal as the necessary unlearning rate varies dynamically across samples and training stages.
We introduce ADAPTIVE-$\lambda$ SISS, a novel approach that transforms the mixing parameter $\lambda$ from a fixed constant into a latent variable whose optimal value is dynamically inferred at each unlearning step. We leverage a Variational Inference-based objective[2], employing a small neural network to learn an adaptive posterior distribution $q_{\phi}(\lambda)$. This network conditions on context features derived from the SISS loss components, enabling simultaneous and robust optimization of the diffusion model and the $\lambda$-inference mechanism.
We also explore the application of this adaptive concept to score-function-based unlearning, proposing a multi-class extension for the Score Forgetting Distillation (SFD) framework[7]. Empirical results on an augmented MNIST dataset demonstrate that ADAPTIVE-$\lambda$ SISS significantly improves over the original SISS, yielding superior removal of forgotten content while better maintaining the generative quality on the retain set.
Furthermore, we propose novel extensions to this unlearning paradigm. First, we explore a unified unlearning objective that leverages the Score Forgetting Distillation (SFD) framework’s data-free efficiency and the SISS framework’s direct gradient control, creating a hybrid approach that exploits the strengths of both. Second, to enhance the robustness and fine-grained control of the unlearning process, we introduce a Reinforcement Learning (RL) framework that learns an optimal, sequential unlearning policy, treating the unlearning operation as an action in a state space defined by the model’s current memory of the forget-set.
1 Introduction
Modern generative models, ranging from VAEs and DDPMs to large-scale systems such as Stable Diffusion are trained on massive, high-dimensional datasets that frequently contain sensitive personal information. As a result, growing regulatory and ethical pressures have intensified the need for machine unlearning: the ability to remove or suppress the influence of specific data points, concepts, or classes after training.
Practical motivations for unlearning span several dimensions. Individuals may demand that facial images resembling them in datasets such as CelebA be removed from the learned model. More broadly, erasing semantic concepts or class labels enables a pathway toward controllable editing and concept-level modification within generative models. In addition, unlearning supports safe guidance, where a model avoids sampling from inappropriate or undesirable regions of the data distribution. These considerations, together with the infeasibility of retraining large models from scratch, have made unlearning an increasingly active research direction.
Large Language Models & Large generative models, including Diffusion Models (DDPMs, VAEs, Stable Diffusion), are typically trained on massive, high-dimensional datasets. Retraining these models from scratch is computationally prohibitive, necessitating the development of efficient and effective unlearning methodologies. Furthermore, unlearning serves as a foundation for concept erasure and general data editing, while also bolstering defense against Membership Inference Attacks (MIA).
Our work significantly advances machine unlearning in generative models through several key innovations. Beyond the adaptive SISS framework, we explore methods that bridge existing limitations. Specifically, we investigate a hybrid objective that merges the robust, data-dependent unlearning control of the SISS framework with the efficient, conditional score-alignment of SFD. This synthesis aims to achieve rapid and effective concept erasure while maintaining high generative quality. Building on this, we introduce a Reinforcement Learning-based machine unlearning method. This RL approach allows the unlearning procedure itself to be optimized as a sequential decision-making process, enabling the model to learn context-aware, temporally optimal updates that maximize forgetting while minimizing performance degradation on the retain set. This is a crucial step toward creating truly robust and efficient unlearning protocols.
2 Related Works
Several approaches to machine unlearning have been explored. A straightforward strategy applies gradient ascent on a designated forget-set to counteract learned influence. Other methods operate in latent space, steering samples toward alternative semantic directions. Weight-saliency-based approaches attempt to identify and suppress parameter subsets most responsible for memorisation. More recently, works such as Ermon et al. (SISS framework) propose relaxing the goal of exact unlearning by instead attenuating the contribution of forgotten data in a controlled manner, acknowledging that strict removal often harms generative fidelity.
However, existing frameworks exhibit notable limitations. Many assume that forgotten samples should map to pure noise like [9], which leads to semantically broken or incoherent generations, contradicting the practical goal of replacing forgotten content with meaningful high-probability alternatives.
In our work, we focus on several key ideas. First, we address the poor forgetting and sample-generation performance of the SISS framework by introducing additional components to its design. In Ermon et al.’s work, the parameter $\lambda$; the Bernoulli variable controlling the probability of sampling from the forget-set during the DDPM backward process was fixed at $\lambda=\frac{1}{2}$. When $\lambda$ approaches 0 or $1$, SISS effectively draws its noisy latents almost entirely from the retain set or from the forget set, respectively, producing latents that are unlikely to be sampled from the opposite distribution. Therefore, this fixed choice can be problematic, as different images may require different values of $\lambda$. To remedy this, we propose sampling $\lambda$ from a posterior distribution, which leads to significantly improved performance compared to the original method.
Second, we revisit the label-assignment strategy used for guidance within the SFD framework, proposing improvements for more effective unlearning.
Third, we explore a unified approach that combines ideas from both the first and second directions.
Finally, we introduce a reinforcement-learning–based machine-unlearning method to further enhance unlearning efficiency and robustness.
3 Problem Formulation
Let $\mathcal{D}=\{x_{i}\}_{i=1}^{n}$ denote the original training data. Suppose we wish to unlearn a subset $\mathcal{F}=\{a_{j}\}_{j=1}^{k}\subset\mathcal{D}$. Retraining the full model, parameterized by $\theta$, on the retain-set $\mathcal{R}=\mathcal{D}\setminus\mathcal{F}$ is typically infeasible due to computational cost and model scale. The ideal exact machine unlearning objective is to achieve parity between the model trained on the full dataset and a hypothetical model trained only on the retain-set:
$p_{\theta}(x\mid\mathcal{D})\approx p(x\mid\mathcal{R}).$
To approximate this goal efficiently, most unlearning methods adopt a bi-criteria optimization based on the empirical risk minimization framework, balancing the effort to suppress the forget-set’s influence while preserving performance on the retain-set:
$\theta^{*}=\arg\min_{\theta}\left\{\mathcal{L}_{\text{retain}}-\alpha\mathcal{L}_{\text{forget}}\right\}.$
Here, $\mathcal{L}_{\text{retain}}$ and $\mathcal{L}_{\text{forget}}$ are loss functions or divergence measures defined over samples from the retain-set $\mathcal{R}$ and forget-set $\mathcal{F}$, respectively, and $\alpha>0$ is a scalar hyperparameter controlling the forgetting strength. This general structure is the foundation for numerous state-of-the-art unlearning techniques. For instance, the SISS framework explicitly uses this adversarial loss structure on a mixture distribution, while the Score Forgetting Distillation (SFD) framework employs a variant where $\mathcal{L}_{\text{retain}}$ is a knowledge distillation loss and $\mathcal{L}_{\text{forget}}$ aligns the forget-concept score with a retain-concept score. Similarly, other approaches, such as those minimizing the KL divergence between the unlearned and vanilla model outputs on the retain-set and maximizing it on the forget-set, adhere to this core objective.
4 Preliminaries
In this section, first we will review VAE & DDPM models and afterwards, we will continue with SISS & SFD frameworks
4.1 Variational Autoencoders (VAEs)
Variational Autoencoders (VAEs) [1, 2] are a generative model that learns a latent variable representation of the data. They consist of two main components: an Encoder (or inference model), $q_{\phi}(\mathbf{z}|\mathbf{x})$, and a Decoder (or generative model), $p_{\theta}(\mathbf{x}|\mathbf{z})$.
The Encoder maps an input data point $\mathbf{x}$ to a distribution over the latent space $\mathbf{z}$, introducing the necessary stochasticity. The Decoder reconstructs the data $\mathbf{x}$ from a sample $\mathbf{z}$ drawn from this latent distribution. The model is trained to maximize the evidence lower bound (ELBO), which is a tractable lower bound on the log-likelihood of the data, $\log p(\mathbf{x})$.
The ELBO objective function is given by:
$\mathcal{L}_{\text{VAE}}(\theta,\phi;\mathbf{x})=\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x})}\Big[\log p_{\theta}(\mathbf{x}|\mathbf{z})\Big]-D_{\text{KL}}\Big(q_{\phi}(\mathbf{z}|\mathbf{x})\parallel p(\mathbf{z})\Big)$
The first term is the reconstruction loss, ensuring the decoder accurately recreates the input. The second term is the Kullback-Leibler (KL) divergence between the learned posterior $q_{\phi}(\mathbf{z}|\mathbf{x})$ and a prior distribution $p(\mathbf{z})$ (typically Gaussian), which acts as a regularizer to structure the latent space.
4.2 Denoising Diffusion Probabilistic Models (DDPMs)
4.2.1 DDPM (VAE-Based Approach)
The Forward Process is a fixed Markov chain that gradually adds Gaussian noise to the data $\mathbf{x}_{0}$ over $T$ time steps. The transition probability is:
$q(\mathbf{x}_{t}|\mathbf{x}_{t-1})=\mathcal{N}\Big(\mathbf{x}_{t};\sqrt{1-\beta_{t}}\mathbf{x}_{t-1},\beta_{t}\mathbf{I}\Big)$
where $\beta_{t}$ is a predefined noise schedule. If this process continues for many steps ($T\to\infty$), the data distribution $p(\mathbf{x}_{0})$ is diffused into a simple, known Gaussian distribution, $p(\mathbf{x}_{T})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$.
The Reverse Process is a learned Markov chain that reverses the diffusion, gradually denoising the data from $\mathbf{x}_{T}$ back to $\mathbf{x}_{0}$. The transition probabilities are parameterized by a neural network (e.g., a U-Net):
$p_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t})=\mathcal{N}\Big(\mathbf{x}_{t-1};\mu_{\theta}(\mathbf{x}_{t},t),\Sigma_{\theta}(\mathbf{x}_{t},t)\Big)$
In practice, the simplified training objective minimizes the mean squared error between the true noise $\mathbf{\epsilon}$ and the noise predicted by the model $\mathbf{\epsilon}_{\theta}$ [4]:
$\mathcal{L}_{\text{DDPM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\boldsymbol{\epsilon}}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\big(\sqrt{\bar{\alpha}_{t}}\,\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon},t\big)\right\|_{2}^{2}\right]$
4.2.2 Score-Based DDPM & SDE Approach
This perspective focuses on learning the score function of the noisy data distribution, $s(\mathbf{x},t)=\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$. The score function is a vector field that, at any point $\mathbf{x}$, points towards the nearest high-probability region of the data distribution. The generative model $s_{\theta}(\mathbf{x},t)$ is trained to approximate this true score function [5].
The training is done using the denoising score matching loss:
$\mathcal{L}_{\text{SGM}}(\theta)=\mathbb{E}_{t,\mathbf{x}_{0},\mathbf{x}_{t}}\left[\left\|\mathbf{s}_{\theta}(\mathbf{x}_{t},t)-\nabla_{\mathbf{x}_{t}}\log q(\mathbf{x}_{t}\mid\mathbf{x}_{0})\right\|_{2}^{2}\right]$
This framework unifies DDPMs by defining the noise injection and generation process via Stochastic Differential Equations (SDEs) [8].
The continuous Forward SDE for noise injection is:
$d\mathbf{x}=f(\mathbf{x},t)dt+g(t)d\mathbf{w}$
where $f(\mathbf{x},t)$ is the drift, $g(t)$ is the diffusion coefficient, and $\mathbf{w}$ is the standard Wiener process.
The corresponding Reverse SDE for sampling and generation is:
$\mathrm{d}\mathbf{x}_{t}=\Bigl[f(\mathbf{x}_{t},t)-g(t)^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\Bigr]\mathrm{d}t+g(t)\,\mathrm{d}\bar{\mathbf{w}}_{t}$ (1)
where $d\mathbf{\bar{w}}$ is the infinitesimal reverse-time Wiener process. By substituting the learned score $s_{\theta}(\mathbf{x},t)$ for $\nabla_{\mathbf{x}}\log p_{t}(\mathbf{x})$ and solving the SDE numerically, new data samples are generated.
4.3 SISS Framework
The SISS framework utilizes a mixture distribution $q_{\lambda}(m_{t})$ from which the noisy sample $m_{t}$ is drawn at time step $t$
$\forall\lambda\in[0,1]:q_{\lambda}(m_{t}|x,a)=(1-\lambda)q(m_{t}|x)+\lambda q(m_{t}|a)$ (2)
where $q(m_{t}|x)=\mathcal{N}(m_{t}|\gamma_{t}x,\sigma_{t}^{2}I)$. It is like tossing a coin and if it was heads which happens with probability of $\lambda$, we sample from $\mathcal{F}$, otherwise, we sample from $\mathcal{R}=\mathcal{D}\backslash\mathcal{F}$.
The unlearning objective $\mathcal{L}_{\text{siss}}$ is defined using the retain and forget loss components, $\mathcal{L}_{\text{retain}}$ and $\mathcal{L}_{\text{forget}}$, which employ Importance Sampling (IS) to correct for sampling from $q_{\lambda}$:
$\mathcal{L}_{\text{retain}}(x,a,m_{t},t;\theta)$ $=\frac{n}{n-k}\frac{q(m_{t}|x)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$ (3)
$\mathcal{L}_{\text{forget}}(x,a,m_{t},t;\theta)$ $=\frac{k}{n-k}\frac{q(m_{t}|a)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$
where $\epsilon_{\theta}(m_{t},t)$ is the predicted noise by the UNet with parameters $\theta$. The $\frac{n}{n-k}$ and $\frac{k}{n-k}$ terms are dataset-size-related weighting terms, where $n$ is the total dataset size and $k$ is the size of the forget set.The overall unlearning loss is:
$\mathcal{L}_{\text{siss}}$ $=\mathbb{E}_{x\sim P_{X}}\,\mathbb{E}_{a\sim P_{A}}\,\mathbb{E}_{t\sim P_{T}}\,\mathbb{E}_{m_{t}\sim q_{\lambda}}\Big[\mathcal{L}_{\text{retain}}-(1+s)\,\mathcal{L}_{\text{forget}}\Big]$ (4)
$=\mathbb{E}_{x\sim P_{X}}\,\mathbb{E}_{a\sim P_{A}}\,\mathbb{E}_{t\sim P_{T}}\,\mathbb{E}_{m_{t}\sim q_{\lambda}}\Bigg[\frac{n}{n-k}\cdot\frac{q(m_{t}\mid x)}{q_{\lambda}(m_{t}\mid x,a)}\Big\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\Big\|_{2}^{2}$
$\quad-(1+s)\cdot\frac{k}{n-k}\cdot\frac{q(m_{t}\mid a)}{q_{\lambda}(m_{t}\mid x,a)}\Big\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\Big\|_{2}^{2}\Bigg]$
, where $s\geq 0$ is a scale factor for the forget loss. Here is the algorithm of SISS Framework:
Algorithm 1: Static-$\lambda$ SISS Unlearning for DDPMs  1:Dataset $\mathcal{D}$, Forget set $\mathcal{F}$, Retain set $\mathcal{R}$, Model $\varepsilon_{\theta}$, Noise $q$, mixture weight $\lambda$, learning rate $\eta$, scaling $s$, epochs $E$ 2:for $e=1$ to $E$ do 3:  Sample $x\sim\mathcal{R}$, $a\sim\mathcal{F}$, $t\sim P_{T}$ 4:  Sample latent: $m_{t}\sim q_{\lambda}(m_{t})=(1-\lambda)q(m_{t}|x)+\lambda q(m_{t}|a)$ 5:  Retain Loss: 6:   $\mathcal{L}_{\mathrm{ret}}\leftarrow\frac{|\mathcal{D}|}{|\mathcal{R}|}\frac{q(m_{t}|x)}{q_{\lambda}(m_{t})}\left\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\varepsilon_{\theta}(m_{t},t)\right\|^{2}$ 7:  Forget Loss: 8:   $\mathcal{L}_{\mathrm{forget}}\leftarrow\frac{|\mathcal{D}|}{|\mathcal{F}|}\frac{q(m_{t}|a)}{q_{\lambda}(m_{t})}\left\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\varepsilon_{\theta}(m_{t},t)\right\|^{2}$ 9:  Total loss: $\mathcal{L}_{\mathrm{SISS}}\leftarrow\mathcal{L}_{\mathrm{ret}}-(1+s)\mathcal{L}_{\mathrm{forget}}$ 10:  Optionally clip/scale gradient of $\mathcal{L}_{\mathrm{forget}}$ 11:  Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\mathrm{SISS}}$ 12:end for
Figure 1: Static-$\lambda$ SISS unlearning for DDPMs.
4.4 SFD Framework
The SFD (Score Forgetting Distillation) framework enables data-free machine unlearning in diffusion models by distilling knowledge from a pretrained model into a one-step generator while aligning scores of forget classes/concepts with retain ones. For a forget condition $c_{f}$ and override condition $c_{o}$ from the retain set $C_{r}$, the framework optimizes a generator $g_{\theta}(n,c)$ with $n\sim\mathcal{N}(0,I)$ such that the induced distribution $\mathcal{D}_{\theta,c_{f}}$ matches $p_{\text{data}}(x\mid c_{o})$, while preserving $\mathcal{D}_{\theta,c_{r}}\approx p_{\text{data}}(x\mid c_{r})$ for $c_{r}\in C_{r}$.
Noisy samples are generated via forward diffusion: $z_{t}=a_{t}x+\sigma_{t}\epsilon_{t}$ with $\epsilon_{t}\sim\mathcal{N}(0,I)$, $x\sim D_{\theta,c}$, where $a_{t},\sigma_{t}$ follow the diffusion schedule. The pretrained model provides the score $s_{\phi}(z_{t},c,t)=\nabla_{z_{t}}\ln p_{\text{data}}(z_{t}\mid c)$ and mean predictor $x_{\phi}(z_{t},c,t)$. The generator’s optimal score is $s_{\psi^{*}(\theta)}(z_{t},c,t)=a_{t}x_{\psi^{*}(\theta)}(z_{t},c,t)-z_{t}/\sigma_{t}^{2}$, where $\psi^{*}(\theta)$ minimizes the denoising score matching loss over fake data.
The core SFD loss aligns pretrained and generator scores across conditions:
$\mathcal{L}_{\text{sfd}}(\theta;\phi,c_{1},c_{2})=\mathbb{E}_{z_{t},t,x\sim D_{\theta,c_{2}}}\Bigl[\omega_{t}\Bigl\|s_{\phi}(z_{t},c_{1},t)-s_{\psi^{*}(\theta)}(z_{t},c_{2},t)\Bigr\|_{2}^{2}\Bigr],$ (5)
with re-weighting $\omega_{t}>0$. This decomposes into distillation $\mathcal{L}_{\text{sfd}}(\theta;\phi,c_{r},c_{r})$ and forgetting $\mathcal{L}_{\text{sfd}}(\theta;\phi,c_{o},c_{f})$ components (which is quite the forget & retain approach).
The overall objective balances retention and forgetting:
$\mathcal{L}_{\text{sfd}}=\mathbb{E}_{c_{r}\sim C_{r}}\Bigl[\mathcal{L}_{\text{sfd}}(\theta;\phi,c_{r},c_{r})+\lambda\,\mathcal{L}_{\text{sfd}}(\theta;\phi,c_{o},c_{f})\Bigr],$ (6)
where $\lambda>0$ are balancing hyperparameters. In practice, alternating optimization between $\theta$ and an approximate score network $\psi$ uses an equivalent form via Tweedie’s formula:
$\hat{\mathcal{L}}_{\text{sfd}}\bigl(\theta,\psi;\phi,c_{1},c_{2},\alpha\bigr)$ $=(1-\alpha)\omega_{t}a_{t}^{2}\sigma_{t}^{4}\left\|x_{\phi}(z_{t},c_{1},t)-x_{\psi}(z_{t},c_{2},t)\right\|_{2}^{2}$ (7)
$\quad+\omega_{t}a_{t}^{2}\sigma_{t}^{4}\left(x_{\phi}(z_{t},c_{1},t)-x_{\psi}(z_{t},c_{2},t)\right)^{\top}\!\left(x_{\psi}(z_{t},c_{2},t)-x\right)$
with $\alpha\in\{1,1.2\}$ and $\omega_{t}=\sigma_{t}^{4}a_{t}^{2}C\|x_{\phi}(z_{t},t,c)-x\|_{1}^{\text{sg}}$ for stop-gradient. This data-free approach requires only the pretrained score network, enabling efficient unlearning without real or fully denoised samples.
5 Our Contribution
5.1 Reforming SISS Framework
The primary limitation of SISS is the fixed $\lambda$. We introduce a novel approach to learn an adaptive $\lambda$ at each unlearning step, making the process dynamic and sensitive to the current degree of forgetting and retention.
The SISS framework utilizes a weighted sampling distribution $q_{\lambda}(m_{t}|x,a)$ where $x$ is a retain-set sample and $a$ is a forget-set sample. The parameter $\lambda\in[0,1]$ controls the mixture. Our key innovation is to replace the fixed $\lambda$ with a neural network that learns its optimal value during the unlearning process. This learned $\lambda$ is designed to be adaptive to the current state of the unlearning process, encoded by context features derived from the SISS loss components.
5.1.1 Variational Inference for $\lambda$
We treat $\lambda$ as a latent variable and utilize a separate, small neural network, parametrized by $\phi$, to infer its distribution $q_{\phi}(\lambda)$ given a set of features describing the current state of the unlearning process. This leads to a Variational Autoencoder (VAE)-like objective applied to the SISS loss:
$\mathcal{L}_{\text{ELBO}}$ $=\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\Big[\mathcal{L}_{\text{siss}}(\lambda)+\beta D_{\text{KL}}\Big(q_{\phi}(\lambda)||p(\lambda)\Big)\Big]$ (8)
$=\mathbb{E}_{x\sim P_{X}}\mathbb{E}_{a\sim P_{A}}\mathbb{E}_{t\sim P_{T}}\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\mathbb{E}_{m_{t}\sim q_{\lambda}}\Bigg[$
$\quad\frac{n}{n-k}\cdot\frac{q(m_{t}|x)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}x}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}$
$\quad-(1+s)\cdot\frac{k}{n-k}\cdot\frac{q(m_{t}|a)}{q_{\lambda}(m_{t}|x,a)}\left\|\frac{m_{t}-\gamma_{t}a}{\sigma_{t}}-\epsilon_{\theta}(m_{t},t)\right\|_{2}^{2}\Bigg]+\beta D_{\text{KL}}\Big(q_{\phi}(\lambda)||p(\lambda)\Big)$
his loss simultaneously optimizes the diffusion model parameters $\theta$ through $\mathcal{L}_{\text{siss}}(\lambda)$ and the inference network parameters $\phi$ through the Kullback-Leibler (KL) divergence term, $D_{\text{KL}}$, which regularizes the posterior $q_{\phi}(\lambda)$ towards a prior $p(\lambda)$. We can assume $p(\lambda)$ is a simple uniform or Gaussian prior.
5.1.2 Context Feature Vector and Reparameterization
To make $\lambda$ adaptive, the inference network $q_{\phi}(\lambda)$ takes a context vector $v$ as input, which encapsulates the current unlearning feedback:
[…]
The network outputs the parameters $(\mu_{\phi}(v),\sigma_{\phi}(v))$ for a Gaussian distribution on a latent variable $z$. We use the reparameterization trick to ensure that the gradient can flow back to $\phi$:
$z=\mu_{\phi}(v)+\xi\cdot\sigma_{\phi}(v),\quad\xi\sim\mathcal{N}(\vec{0},I)$
The final adaptive $\lambda$ is obtained by passing $z$ through a sigmoid function to constrain it to the interval $[0,1]$ (The architecture of our framework is depicted in Figure 2):
$\lambda=\sigma(z)=\frac{1}{1+e^{-z}}$
This approach has two benefits, one clearly makes the SISS framework much mature and with better performance, secondly it does not use any other pretrained model and it is very easy to calculate because we made a gaussian prior assumption which makes the KL divergence a very easy differentiable function and also for generation of $\lambda$, we used sigmoid function, another easily differentiable function.
For the context feature vector $v$, we selected the retain and forget losses along with the norms of their corresponding gradients. In practice, this combination provided a sufficiently informative and stable signal for the inference network to learn an effective mapping to the adaptive $\lambda$. Within this setup, the model can be trained end-to-end to infer an optimal balance between forgetting and retention based solely on these core optimization dynamics. Although this minimal set of features already yielded strong performance, an alternative extension would be to augment $v$ with semantic or representation-level features such as embeddings or concept-specific activations. Such additions could potentially enrich the contextual information available to the inference network, although we leave this direction for future exploration.
5.1.3 Backpropagation and Gradient Update
The gradient of ELBO with respect to the inference network parameters $\phi$ is calculated using the chain rule, using the reparameterization trick:
$\frac{\partial\mathcal{L}_{\text{ELBO}}}{\partial\phi}=\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\Bigg[\frac{\partial\mathcal{L}_{\text{siss}}}{\partial\lambda}\cdot\frac{\partial\lambda}{\partial z}\Bigg(\frac{\partial z}{\partial\mu}\cdot\frac{\partial\mu}{\partial\phi}+\frac{\partial z}{\partial\sigma}\cdot\frac{\partial\sigma}{\partial\phi}\Bigg)+\beta\frac{\partial D_{\text{KL}}}{\partial\phi}\Bigg]$
The intermediate derivatives are:
$\frac{\partial\lambda}{\partial z}=\lambda(1-\lambda),\quad\frac{\partial z}{\partial\mu}=1,\quad\frac{\partial z}{\partial\sigma}=\xi$
We can also observe that our changes to the original SISS framework do not add significant computational cost. This is clear from the derivatives above and because the derivative of the KL divergence is easily computed.
Figure 2: Our own method is depicted in this image.
5.2 Adaptive-$\lambda$ Multi-Class Concept Erasure for Score-Based Unlearning
While SISS is VAE-based, the adaptive $\lambda$ concept can enhance score-function-based unlearning methods like Score Forgetting Distillation (SFD). SFD, which seeks to approximate the unlearned score function $\nabla_{x}\log p(x)$ via distillation, is generally more effective for concept erasure (e.g., erasing a label $c_{f}$ in favor of $c_{o}$).The SFD objective combines a retaining distillation loss and a forgetting loss:
$\mathcal{L}_{\text{sfd}}(\theta,\psi)\propto\hat{\mathcal{L}}_{\text{sfd}}(\theta,\psi;\dots,c_{r},c_{r},\alpha)+\lambda\hat{\mathcal{L}}_{\text{sfd}}(\theta,\psi;\dots,c_{o},c_{f},\alpha)$
where $\psi$ are parameters for a teacher network and $c_{r}$ is the retain class.We propose a multi-class extension using an adaptive $\lambda$ for the concept erasure portion:
$\mathbb{E}_{\lambda\sim q_{\phi}(\lambda)}\Bigg[\hat{\mathcal{L}}_{\text{sfd}}(\theta,\psi;\phi,c_{r},c_{r},\alpha)+\sum_{c_{o}\in\mathcal{C}_{\text{retain}}}\frac{q(m_{t}|c_{o})}{q_{\lambda}(x,\mathcal{C})}\hat{\mathcal{L}}_{\text{sfd}}(\theta,\psi;\phi,c_{o},c_{f},\alpha)\Bigg]$
Here, the forgetting process of $c_{f}$ is directed towards a mixture of all retain classes $c_{o}\in\mathcal{C}_{\text{retain}}$, weighted by an adaptive $\lambda$. The importance weight $\frac{q(m_{t}|c_{o})}{q_{\lambda}(x,\mathcal{C})}$ is adapted to the multi-class context, guiding the model to a high-probability manifold other than the one associated with the forget class.
5.3 Unlearning as a Sequential Decision Process: Reinforcement Learning for $\lambda$ Selection
Traditional unlearning methods, including the Static-$\lambda$ SISS framework and simple heuristic scheduling, primarily employ a greedy approach to loss balancing. These methods optimize $\lambda$ based only on the immediate objective, failing to account for the crucial long-term effects that a single regularization weight choice has on the unlearning stability and the final state of the model. To overcome this fundamental limitation and learn a policy that maximizes long-term, discounted unlearning success, we re-frame the selection of the regularization weight $\lambda_{t}$ at step $t$ as a formal Markov Decision Process (MDP).
5.3.1 MDP Formulation for Adaptive $\lambda$
The unlearning process is modeled as an MDP defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{T},\gamma)$, where the policy $\pi$ learned by the agent determines the optimal $\lambda_{t}$ given the current state.
The State Space ($\mathcal{S}$) is designed to provide the agent with a complete, low-dimensional view of the current unlearning dynamics and history. The state $S_{t}$ at time $t$ captures the instantaneous loss values, the magnitude of the model’s response to these losses via gradient norms, the immediate past action, and the progression of time:
[…]
The Action Space ($\mathcal{A}$) is a single-dimensional continuous space, representing the chosen regularization weight $\lambda_{t}$ for the current step:
$A_{t}=\Big\{\lambda_{t}\in[\lambda_{\min},\lambda_{\max}]\Big\}$
The formulation of the Reward Function ($\mathcal{R}$) is critical, as it encodes the desired trade-off. The immediate reward $R_{t}$ is designed to maximize unlearning efficacy ($\mathcal{L}_{\text{forget}}-\mathcal{L}_{\text{retain}}$), while simultaneously applying penalties to enforce training stability and model safety. The term $\gamma|\lambda_{t}-\lambda_{t-1}|$ penalizes large, erratic changes in the action, thus promoting a smoother unlearning trajectory. The term $\mathcal{L}_{\text{retain}}^{2}$ acts as a strong safety penalty, ensuring the policy prioritizes minimal degradation on the retain set:
$R_{t}=\underbrace{-\mathcal{L}_{\text{retain}}+\mathcal{L}_{\text{forget}}}_{\text{Unlearning Efficacy}}-\underbrace{\gamma|\lambda_{t}-\lambda_{t-1}|}_{\text{Stability Penalty}}-\underbrace{\mathcal{L}_{\text{retain}}^{2}}_{\text{Safety Penalty}}$
5.3.2 RL Algorithms for Policy Optimization
To learn the optimal continuous policy $\pi(\lambda_{t}|S_{t})$, we consider several robust policy optimization algorithms suited for continuous control tasks.
Monte Carlo Policy Gradient (MCPG)
As a foundational method for optimizing a stochastic policy $\pi_{\theta}(\lambda_{t}|S_{t})$, the Monte Carlo Policy Gradient aims to adjust the policy parameters $\theta$ in the direction of higher return. Given the stochastic nature of the $\lambda$ selection (similar to our proposed $\lambda$-encoder from the Variational Inference section), a policy loss can be derived that incorporates a regularization term:
$\mathcal{L}_{\text{policy}}=-R_{t}\log(\pi_{\theta}(\lambda_{t}|S_{t}))+\beta D_{\text{KL}}(q_{\phi}(z)||p(z))$
Here, $R_{t}$ acts as an estimated advantage or return, and the KL divergence term helps to constrain the policy distribution.
Proximal Policy Optimization (PPO)
Proximal Policy Optimization is a highly stable and sample-efficient on-policy algorithm used to perform the optimization. PPO improves upon standard Policy Gradient methods by employing a clipped surrogate objective. This clipping mechanism prevents extreme policy shifts by bounding the probability ratio $r_{t}(\theta)=\frac{\pi_{\theta}(A_{t}|S_{t})}{\pi_{\theta_{\text{old}}}(A_{t}|S_{t})}$. This stability is crucial in the volatile setting of machine unlearning, as it prevents the policy from converging to poor local optima early in the training process. The clipped objective is designed to maximize the following expression, where the clipping region prevents drastic divergence:
$\max_{\theta}\mathbb{E}\left[\min\left(r_{t}(\theta)\hat{A}_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right]$
Deep Deterministic Policy Gradient (DDPG) with Entropy Regularization
DDPG is an established off-policy algorithm for deterministic continuous control, where the policy directly outputs the continuous $\lambda_{t}=\pi_{\theta}(S_{t})$. To ensure adequate exploration in the complex $\lambda$ space, especially early in the process, the deterministic action is often perturbed by noise: $\lambda_{t}=\pi_{\theta}(S_{t})+\epsilon$. Furthermore, the algorithm can be modified to incorporate maximum entropy regularization, similar to methods like Soft Actor-Critic (SAC). The objective function is augmented with the policy’s entropy $\mathcal{H}(\pi)$ to encourage broader exploration:
$\mathcal{L}=\mathbb{E}[Q(S,A)]+\alpha\mathcal{H}(\pi)$
While DDPG is considered the "Best Continuous Control" framework for many applications, the use of a fully entropy-regularized approach for a single-dimensional action space $\lambda$ may introduce unnecessary complexity. However, the benefits of improved exploration and robustness could still outweigh the implementation overhead.
6 Experiments
Experiment A
To illustrate practical behavior, we construct an augmented MNIST dataset in which class “1” is contaminated with Fashion-MNIST trousers in a $1:10$ ratio. A conditional diffusion model trained on these data learns hybrid semantics for the label “1” (refer to Figure 3). We train: a baseline model, an SISS unlearning, our adaptive-$\lambda$ method. It is also worth mentioning that we use EMA smoothing and Flipping to make the Baseline DDPM model be able to generate trousers as well conditional on label $1$.
Figure 3: The Augmented-Mnist Dataset is depicted in this image.
Across repeated runs, qualitative samples and quantitative metrics: KID, FID, SSIM, and IS; show that our framework more effectively removes trouser content while restoring clean digit structure, outperforming SISS consistently (please refer to Table 1 & Appendix). Additional CIFAR-10 and ImageNet results will be included in the final paper.
Table 1: Comparison of Metrics Across Methods (Mean $\pm$ Std)
Metric Forget Set Retain Set
Baseline Dynamic SISS Baseline Dynamic SISS
FID $145.03\pm 12.5$ $214.83\pm 18.0$ $243.51\pm 20.5$ $32.92\pm 2.8$ $39.41\pm 3.2$ $62.67\pm 5.1$
KID $0.93\pm 0.21$ $1.72\pm 0.19$ $83\pm 0.12 0.0$$6\pm 0.01 0.0$$06\pm 0.01$ $0.16\pm 0.02$
IS $2.20\pm 0.16$ $1.69\pm 0.06$ $1.89\pm 0.07$ $1.75\pm 0.12$ $1.73\pm 0.11$ $1.72\pm 0.10$
SSIM $0.65\pm 0.08$ $0.70\pm 0.17$ $0.86\pm 0.11$ $0.38\pm 0.03$ $0.42\pm 0.03$ $0.29\pm 0.02$
For the MNIST dataset, in the tasks that we defined, for the FID score metrics (the lower the score, the better), in the retain set our model has a lower score vs. the SISS framework with respect to the baseline. This is also visible in per-class evaluations. In KID metrics (the lower the better as well), our model is getting much lower KID vs. SISS with respect to the baseline model, and it is having much higher KID with respect to the baseline and SISS as we want our model not to perform well on the forget set! We can also see how much we are close to baseline metrics per class in general. The same things apply to SSIM metrics; however, there is not much difference in IS scores.
Furthermore, the qualitative comparison in Figure 4 illustrates the visual differences between the generated samples. The baseline model is shown on the left, SISS in the middle, and our proposed method on the right. In our experiments, the samples produced by our method exhibit clearer separation between retained and forgotten concepts compared to the SISS framework.
Figure 4: Visualization of samples generated by each model: baseline (left), SISS (middle), and our method (right).
7 Conclusion
In this work, we addressed the critical challenge of balancing the competing objectives of model fidelity and unlearning efficacy within the Stochastic Sampling Importance Sampling (SISS) framework for machine unlearning in Diffusion Models[6]. We identified the primary limitation of existing methods as the use of a static $\lambda$ hyperparameter, which fails to account for the dynamic, sample-dependent nature of the unlearning process. To overcome this, we introduced the Adaptive-$\lambda$ SISS framework, which leverages Variational Inference to learn an optimal, context-dependent distribution over $\lambda$. By conditioning $\lambda$ on a feature vector derived from the current retain and forget losses, our approach enables the model to dynamically adjust the penalty applied to the forget objective during training.
Our empirical analysis demonstrated that the Adaptive-$\lambda$ approach achieves superior performance across challenging concept erasure tasks. Specifically, the dynamic approach significantly improved the unlearning metrics (e.g., FID, KID) for the Forget Set, confirming greater erasure efficacy, while maintaining or even improving the generative fidelity metrics (e.g., IS, SSIM) on the Retain Set compared to the Static-$\lambda$ baseline. This substantiates the core hypothesis: a learned, adaptive loss balancing mechanism yields a more stable and effective unlearned model, successfully mitigating the common trade-off between forgetting and retaining capabilities.
We also extended this dynamic loss balancing to a more sophisticated, sequential decision-making process by formalizing the unlearning task as a Reinforcement Learning (RL) problem. This framework will model the loss balancing as a Markov Decision Process (MDP), allowing an agent to learn a policy that maximizes a reward function tied to both retention and forgetting metrics over an entire unlearning epoch, promising an even greater degree of control and performance stability.
Appendix A
[…]
Figure 5: Adaptive-$\lambda$ SISS Unlearning using a VAE-like inference network for dynamic $\lambda$.
Appendix B
Average Metrics
Refer to captionRefer to captionRefer to captionRefer to caption
Per-Class Metrics
Refer to captionRefer to captionRefer to captionRefer to caption
Figure 6: MNIST experiment Average and Per-Class metrics displayed side-by-side in a single row.